OpenStudy (anonymous):

Find the limit of this: lim n->infinity of (log(n+1) + a^(n+1)) / (log(n) + a^n)

OpenStudy (anonymous):

after doing l'hopital's rule once, i got \[\lim_{n \rightarrow \infty} \frac{ 1/(n+1) + (n+1)a^n }{ 1/n + n * a^{n-1} }\] any idea how to proceed after this?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

A little algebra. \(\dfrac{n\cdot\left(1+(n+1)^2\cdot a^{n}\right)}{(n+1)\cdot\left(1+n^2\cdot a^{n+1}\right)}\) What is the value of 'a'? Any restrictions? You do know that l'Hospital doesn't always work, right? I'm not saying it doesn't here. Just mentioning it.

OpenStudy (anonymous):

Notice that the \(\dfrac{1}{n+1}\) and \(\dfrac{1}{n}\) terms disappear as \(n\to\infty\). Meanwhile, \[\frac{(n+1)a^n}{na^{n-1}}=a\frac{n+1}{n}\]